OpenStudy (anonymous):

Five members of the soccer team and five members of the track team ran the 100-meter dash. Their times are listed in the table below: Soccer Track 12.3 12.3 13.2 11.2 12.5 11.7 11.3 12.2 14.4 13.7 What is the difference of the means for the two groups? 0.52 12.22 12.74 24.96 @ganeshie8

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@iambatman @Directrix

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

Add up the five times of the Soccer Team and divide by 5.

OpenStudy (anonymous):

i got 12.74

Directrix (directrix):

While I check that, add the 5 times of the track team and divide by 5.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got 12.22

OpenStudy (anonymous):

then i added them both together and got 24.96 as the answer?

Directrix (directrix):

I agree with the Soccer team. (12.3 + 13.2 + 12.5 + 11.3 + 14.4)/5= 12.74

Directrix (directrix):

>>What is the difference of the means for the two groups? Don't add the means. Subtract.

Directrix (directrix):

I'll check the track average while you subtract.

OpenStudy (anonymous):

ohhh ok so that qould equal 0.52 :)

Directrix (directrix):

(12.3 + 11.2+ 11.7 + 12.2 + 13.7)/5= 12.22

Directrix (directrix):

12.74 - 12.22 =

Directrix (directrix):

.52 which is your answer.

OpenStudy (anonymous):

thank you soo much i've been stuck on this for 2 days and everyone i'd tag wouldn't even help me i very much appreciate your help thank you so much :)

Directrix (directrix):

You are welcome. You did all the work. It's just that addition rather than subtraction part that messed you up. It is great that you did not give up.

OpenStudy (anonymous):

thx :) @Directrix